OpenStudy (anonymous):

write another fraction that names each of these fractions- three-fourths, two-thirds, two-sixths?

OpenStudy (anonymous):

yes it's my daughters homework I'm rusty on equivalent fractions

OpenStudy (anonymous):

three-fourths 6/6, two-thirds 4/9, two-sixths 4/12

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

still not sure how you got the answer?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Take the numerator and denominator of the beginning fraction and multiply both by the same number. It is the equivalent of multiplying by one. It does not change the value of the fraction, just the terms of the fraction. Example\[\frac{3}{4}=\frac{3}{4}\frac{2}{2}=\frac{6}{8}\]

OpenStudy (anonymous):

myko has a typo in the first line. The other answers are correct.

OpenStudy (anonymous):

do I choose any number? wouldn't 2/3= 4/6 not 4/9